Two traveling sinusoidal waves are described by eq
Answer #1

y_{1}= 3.00[\sin (2.00x-1500t)]

and   y_{2}= 2.00[\sin (2.00x-1500t-0.500)]

y= y_{1}+y_{2} = A'\sin (2.00x-\theta )

Where A' = \sqrt{3^{2}+2^{2}+2\times 3\times 2\cos (0.500)} = \sqrt{3^{2}+2^{2}+2\times 3\times 2\times 0.8775}

A' = 4.85 cm

Part B frequency of the resultant wave function will be zero